# Beyond boundaries: Gary Lorden’s groundbreaking contributions to sequential analysis

### v1

@@ 2.5 Lorden’s 2-SPRT and the Kiefer–Weiss Minimax Optimality
It turns out that it is possible to determine the nearly least favorable point $\theta_{\star}(\alpha_{0},\alpha_{1},\theta_{0},\theta_{1})$ such that the $2$-SPRT with thresholds $a_{i}=\log(1/\alpha_{i})$ is second-order asymptotically minimax for the exponential family (37), i.e., the residual term in the difference between the expectation of the sample size of the optimal test and the 2-SPRT is of order $O(1)$ for small error probabilities. This problem was first addressed by Huffman (1983) who suggested the $\theta_{\star}(\alpha_{0},\alpha_{1},\theta_{0},\theta_{1})$ which leads to the residual term of order $o(|\log{\alpha_{\rm max}}|^{1/2})$. This result was strengthened by Dragalin and Novikov (1987) who have established the second-order optimality of Huffman’s version of 2-SPRT. Specifically, writing
$S_{n}^{\theta}=S_{n}-{\mathsf{E}}_{\theta}[X_{1}]n\equiv S_{n}-\overset{\bm{.}}{b}(\theta)n$
and noting that
$\lambda_{i}^{\theta}(n):=\sum_{t=1}^{n}\log\left[\frac{f_{\theta}(X_{t})}{f_{\theta_{i}}(X_{t})}\right]=(\theta-\theta_{i})S_{n}^{\theta}+I(\theta,\theta_{i})n,$
it is easy to see that the stopping times $T_{i}^{\theta}$ can be written as
$T_{0}^{\theta}$ $=\inf\left\{n:\lambda_{1}^{\theta}(n)\geqslant|\log\alpha_{1}|\right\}=\inf\left\{n:S_{n}^{\theta}\leqslant(\theta_{1}-\theta)^{-1}[-|\log\alpha_{1}|+I(\theta,\theta_{1})n]\right\},$
$T_{1}^{\theta}$ $=\inf\left\{n:\lambda_{0}^{\theta}(n)\geqslant|\log\alpha_{0}|\right\}=\inf\left\{n:S_{n}^{\theta}\geqslant(\theta-\theta_{0})^{-1}[|\log\alpha_{0}|-I(\theta,\theta_{0})n]\right\}.$
Therefore, the stopping time $T^{\star}(\theta)=\min(T_{0}^{\theta},T_{1}^{\theta})$ of the $2$-SPRT tuned to $\theta$ can be represented as
$T^{\star}(\theta)=\min\left\{n\geqslant 1:S_{n}^{\theta}\notin(h_{0}^{\theta}(n),h_{1}^{\theta}(n))\right\},$
where the boundaries $h_{1}^{\theta}(n)$ and $h_{0}^{\theta}(n)$ are linear functions of $n$:
$h_{1}^{\theta}(n)=\frac{|\log\alpha_{0}|-I(\theta,\theta_{0})\,n}{\theta-\theta_{0}},\quad h_{0}^{\theta}(n)=\frac{-|\log\alpha_{1}|+I(\theta,\theta_{1})\,n}{\theta_{1}-\theta}.$
Define $\gamma(\theta,\theta_{i})=(\theta-\theta_{i})/I(\theta,\theta_{i})$. Since $\gamma(\theta,\theta_{1})<0<\gamma(\theta,\theta_{0})$, it is easily verified that these boundaries intersect at the point
$n^{\star}(\theta)=\frac{1}{\gamma(\theta,\theta_{0})+|\gamma(\theta,\theta_{1})|}\left[\frac{\gamma(\theta,\theta_{0})}{I(\theta,\theta_{1})}|\log\alpha_{1}|+\frac{|\gamma(\theta,\theta_{1})|}{I(\theta,\theta_{0})}|\log\alpha_{0}|\right].$ (38)
Therefore, the region of continuation of observations is a triangle, as shown in Figure 1. This means that the $2$-SPRT is a truncated test with the maximal number of steps $n^{\star}(\theta)$. Figure 1 also shows the typical curved boundaries of the optimal truncated test that can be computed using Bellman’s backward induction.
Figure 1: The boundaries $h_{1}^{\theta}(n)$ and $h_{0}^{\theta}(n)$ of the $2$-SPRT (solid) and optimal boundaries (dashed) as functions of $n$.
By Corollary 5.3.1 in Tartakovsky, Nikiforov, and Basseville (2015), the $2$-SPRT $\delta^{\star}(\theta)$ minimizes all positive moments of the stopping time distribution at the point $\theta$ to first order (as $\alpha_{\max}\to 0$) in the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ as long as the ratio $|\log\alpha_{0}|/|\log\alpha_{1}|$ is bounded away from zero and infinity,
$\lim_{\alpha_{\max}\to 0}\frac{|\log\alpha_{0}|}{|\log\alpha_{1}|}=c,\quad 0<c<\infty,$ (39)
and the following asymptotics hold:
$\inf_{\delta\in{\mathbb{C}}(\alpha_{0},\alpha_{1})}{\mathsf{E}}_{\theta}[T]\sim{\mathsf{E}}_{\theta}[T^{\star}(\theta)]\sim\min\left[\frac{|\log\alpha_{1}|}{I(\theta,\theta_{0})},\frac{|\log\alpha_{0}|}{I(\theta,\theta_{1})}\right]\quad\text{as}~{}~{}\alpha_{\max}\to 0.$ (40)
The right-hand side of the asymptotic equality (40) is maximized at $\theta=\theta^{*}$ that satisfies the equation
$\frac{|\log\alpha_{1}|}{I(\theta^{*},\theta_{0})}=\frac{|\log\alpha_{0}|}{I(\theta^{*},\theta_{1})}(\equiv n^{\star}(\theta^{*}))$ (41)
in which case
$\inf_{\delta\in{\mathbb{C}}(\alpha_{0},\alpha_{1})}{\mathsf{E}}_{\theta^{*}}[T]\sim{\mathsf{E}}_{\theta^{*}}[T^{\star}(\theta^{*})]\sim n^{\star}(\theta^{*})\quad\text{as}~{}~{}\alpha_{\max}\to 0$
and $\sup_{\theta}{\mathsf{E}}_{\theta}[T^{\star}(\theta^{*})]\approx n^{\star}(\theta^{*})$. Next, we observe that the function $n^{\star}(\theta)$ attains its minimum at the point $\theta=\theta^{*}$. Thus, the point $\theta^{*}$ has two properties simultaneously – it minimizes the time of truncation $\min_{\theta\in(\theta_{0},\theta_{1})}n^{\star}(\theta)=n^{\star}(\theta^{*})$ on one hand, and approximately maximizes the expected sample size of the $2$-SPRT on the other hand. Hence, we may expect that $\delta^{\star}(\theta^{*})$ is asymptotically minimax. This is indeed the case
$\inf_{\delta\in{\mathbb{C}}(\alpha_{0},\alpha_{1})}~{}\sup_{\theta\in\Theta}{\mathsf{E}}_{\theta}[T]\sim\sup_{\theta\in\Theta}{\mathsf{E}}_{\theta}[T^{\star}(\theta^{*})]\sim{\mathsf{E}}_{\theta^{*}}[T^{\star}(\theta^{*})]\sim n^{\star}(\theta^{*}),$
so that the $2$-SPRT tuned to the point $\theta^{*}$, which is found from (41), is first-order minimax. However, with this choice the supremum of ${\mathsf{E}}_{\theta}[T^{\star}(\theta^{*})]$ over $\theta$ is attained at the point $\theta^{*}$ only to within the term $o(n^{\star}(\theta^{*}))$. This result can be improved to within $O(1)$ by choosing the tuning point in a slightly different manner as $\theta_{\star}=\theta^{*}+C\cdot n^{\star}(\theta^{*})^{-1/2}$. Specifically, Huffman (1983) suggested to use the tuning point
$\theta_{\star}=\theta^{*}+\frac{r^{*}}{\sigma_{\theta^{*}}\sqrt{n^{\star}(\theta^{*})}},$ (42)
where $r^{*}$ is a solution of the equation
$\Phi(r^{*})=\frac{|\gamma(\theta^{*},\theta_{1})|}{|\gamma(\theta^{*},\theta_{1})|+\gamma(\theta^{*},\theta_{0})}$
($\Phi(x)$ is the cdf of the standard normal variable). The residual term in Huffman’s approximation is of order $o(n^{\star}(\theta^{\star})^{1/2})$. Dragalin and Novikov (1987) found a more accurate estimate of the residual term in the expansion for the expected sample size showing that it is of order $O(1)$.
The following theorem states the final result. We recall that $n^{\star}(\theta)$ is defined in (38) and, as usual, we write $\varphi(x)={\mathrm{d}}\Phi(x)/{\mathrm{d}}x$ for density of the standard normal distribution $\Phi(x)$.
Theorem 5.
Let the observed random variables $X_{1},X_{2},\dots$ be i.i.d. from the exponential family (37). Let the thresholds in the $2$-SPRT be selected as $a_{1}=|\log\alpha_{1}|$ and $a_{0}=|\log\alpha_{0}|$, and let the tuning point $\theta_{\star}$ be as in (42). If the condition (39) holds, then as $\alpha_{\max}\to 0$,
${\mathsf{E}}_{\theta_{\star}}[T^{\star}(\theta_{\star})]$ $=n^{\star}(\theta^{*})-\sigma_{\theta^{*}}\left[(\gamma(\theta^{*},\theta_{0})+|\gamma(\theta^{*},\theta_{1})|)\varphi(r^{*})\sqrt{n^{\star}(\theta^{*})}\right]+O(1),$
$\sup_{\theta\in\Theta}{\mathsf{E}}_{\theta}[T^{\star}(\theta_{\star})]$ $={\mathsf{E}}_{\theta_{\star}}[T^{\star}(\theta_{\star})]+O(1),$
$\inf_{\delta\in{\mathbb{C}}(\alpha_{0},\alpha_{1})}\sup_{\theta\in\Theta}{\mathsf{E}}_{\theta}[T]$ $={\mathsf{E}}_{\theta_{\star}}[T^{\star}(\theta_{\star})]+O(1).$
Therefore, the $2$-SPRT $\delta^{\star}(\theta_{\star})$ is second-order asymptotically minimax.
@@
As we have already pointed out, the formulas $a_{i}=|\log\alpha_{i}|$, which guarantee the inequalities $\alpha_{i}(\delta^{\star}(\theta))\leqslant\alpha_{i}$, are rather conservative. An improvement can be obtained by observing that
@@
and by noticing that asymptotically as $a_{i}\to\infty$,
${\mathsf{P}}_{\theta}(T^{\star}=T_{1}^{\theta})\to\frac{|\gamma(\theta,\theta_{1})|}{|\gamma(\theta,\theta_{1})|+\gamma(\theta,\theta_{0})},\quad{\mathsf{P}}_{\theta}(T^{\star}=T_{0}^{\theta})\to\frac{\gamma(\theta,\theta_{0})}{|\gamma(\theta,\theta_{1})|+\gamma(\theta,\theta_{0})},$
@@
For the nonarithmetic case $\zeta_{i}^{\theta}$ can be computed using the renewal-theoretic argument similar to (17) and (19):
$\zeta_{i}^{\theta}=\frac{1}{I(\theta,\theta_{i})}\exp\left\{-\sum_{n=1}^{\infty}\frac{1}{n}\left[{\mathsf{P}}_{\theta}(\lambda_{i}^{\theta}(n)>0)+{\mathsf{P}}_{i}(\lambda_{i}^{\theta}(n)\leqslant 0)\right]\right\}.$
This yields
$\alpha_{0}(\delta^{\star}(\theta))\sim\frac{|\gamma(\theta,\theta_{1})|}{|\gamma(\theta,\theta_{1})|+\gamma(\theta,\theta_{0})}\zeta_{0}^{\theta}e^{-a_{0}},\quad\alpha_{1}(\delta^{\star}(\theta))\sim\frac{\gamma(\theta,\theta_{0})}{|\gamma(\theta,\theta_{1})|+\gamma(\theta,\theta_{0})}\zeta_{1}^{\theta}e^{-a_{1}}.$
@@ 2.6 Near Uniform Optimality of the GLR SPRT for Composite Hypotheses
For practical purposes, it is much more important to design tests that would minimize the expected sample size ${\mathsf{E}}_{\theta}[T]$ for all possible parameter values (i.e., uniformly optimal) rather than to solve the minimax Kiefer–Weiss problem of minimizing ${\mathsf{E}}_{\theta}[T]$ at a least favorable point. In this section, our ultimate goal is to discuss how to design sequential tests that are at least approximately uniformly optimal for small error probabilities or asymptotically Bayesian for small cost of observations for testing composite hypotheses. Assume that the sequence of i.i.d. observations $X_{1},X_{2},\dots$ comes from a common distribution ${\mathsf{P}}_{\theta}$ with density $f_{\theta}(x)$ with respect to some non-degenerate sigma-finite measure, where the $\ell$-dimensional parameter $\theta=(\theta_{1},\dots,\theta_{\ell})$ belongs to a subset $\Theta$ of the Euclidean space $\mathbb{R}^{\ell}$. The parameter space $\Theta$ is split into $3$ disjoint sets $\Theta_{0},\Theta_{1}$ and ${\mathsf{I}_{\mathrm{in}}}$, i.e., $\Theta=\Theta_{0}\,{\textstyle\bigcup}\,\Theta_{2}\,{\textstyle\bigcup}\,{\mathsf{I}_{\mathrm{in}}}$. The problem is to test the two composite hypotheses $\operatorname{\mathcal{H}}_{0}:\theta\in\Theta_{0}$ against $\operatorname{\mathcal{H}}_{1}:\theta\in\Theta_{1}$. The subset ${\mathsf{I}_{\mathrm{in}}}$ of $\Theta$ represents an indifference zone where the loss $L(\theta,d)$ associated with correct or incorrect decisions $d$ is zero, i.e., no constraints on the probabilities ${\mathsf{P}}_{\theta}(d=i)$ are imposed if $\theta\in{\mathsf{I}_{\mathrm{in}}}$. The indifference zone is usually introduced since, in most applications, the correct action is not critical and often not even possible when the hypotheses are very close. However, in principle ${\mathsf{I}_{\mathrm{in}}}$ may be an empty set.
We are interested in finding a sequential test $\delta=(T,d)$ that minimizes the expected sample size ${\mathsf{E}}_{\theta}[T]$ uniformly for all $\theta\in\Theta$ in the class of tests ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ in which the maximal error probabilities $\sup_{\theta\in\Theta_{i}}{\mathsf{P}}_{\theta}(d\neq i)$ are upper-bounded by the given values:
@@
In other words, we are interested in the following frequentist problem: Find a test $\delta_{\rm opt}$ such that
@@
Kiefer and Sacks (1963) showed that the procedure $\widetilde{\delta}(c)=(\widetilde{T}(c),\widetilde{d}(c))$ with the stopping time
$\widetilde{T}(c)=\inf\left\{n\geqslant 1:R_{n}^{\rm st}(S_{n})\leqslant c\right\}$
proposed by Schwarz (1962), is also first-order asymptotically Bayes, i.e., for any prior distribution $W$, $\rho_{c}^{W}(\widetilde{\delta}(c))\sim\inf_{\delta}\rho_{c}^{W}(\delta)$ as $c\to 0$. Lorden (1967) refined this result by introducing the stopping region as the first $n$ such that $R_{n}^{\rm st}(S_{n})\leqslant Qc$, where $Q$ is a positive constant, and showed that it can be made second-order asymptotically optimal, i.e., $\inf_{\delta}\rho_{c}^{W}(\delta)=\rho_{c}^{W}(\widetilde{\delta}(Qc))+O(c)$ as $c\to 0$, while $\inf_{\delta}\rho_{c}^{W}(\delta)=O(c|\log c|)$. Notice that the problem addressed by Lorden (1967) is much more general than what we are discussing here, since it encompasses general i.i.d. models – not only exponential families – and multiple-decision cases.
A breakthrough in the Bayesian theory for testing separated hypotheses about the parameter of the one-parameter exponential family (37) was achieved by Lorden (1977b) (an unpublished manuscript), where he showed that the family of GSLRTs can be designed to guarantee the third-order asymptotic optimality, i.e., to attain the Bayes risk to within $o(c)$ as $c\to 0$. Lorden gave sufficient conditions for families of tests to be third-order asymptotically Bayes and examples of such procedures based not only on the GLR approach but also on mixtures of likelihood ratios. In addition, the error probabilities of the GSLRTs have been evaluated asymptotically as a consequence of a general theorem on boundary-crossing probabilities. Due to the importance of this work, we give a more detailed overview of Lorden’s theory. Notice that the paper Lorden (1977b) is an extension to a continuous parameter case of the results obtained by Lorden (1977a) for multiple discrete cases that we discussed in Subsection 2.4.
@@
for the LLR between points $\theta$ and $\theta_{i}$.
@@
Note the crucial difference between Schwarz’s GSLRT (47) and Lorden’s GSLRT (50). While in the Schwarz test $h_{i}\equiv 1$ and the threshold is $a=|\log c|$, in the Lorden test, there are two innovations. First, the threshold is smaller by $\tfrac{1}{2}\log|\log c|$, and second, there are adaptive weights $h_{i}(\hat{\theta}_{n})$ in the GLR statistic. Since the stopping times $\widehat{T}_{i}$ can be obviously written as
@@
Lorden’s GSLRT can alternatively be viewed as the GSLRT with curved adaptive boundaries
@@
that depend on the behavior of the MLE $\hat{\theta}_{n}$. These two innovations make this modification of the GLR test nearly optimal.
The formal mathematical proof by Lorden is very involved, so we give only a heuristic sketch that fixes the main ideas of the approach. The Bayesian consideration naturally leads to the mixture LR statistics
@@
where $i=0$ if $\hat{\theta}_{n}\leqslant\theta^{*}$ and $i=1$ otherwise. These approximations follow from the well-known asymptotic expansions for integrals, i.e., Laplace’s integration method and its variations.
Next, Lorden showed that there exists $Q>1$ such that if the stopping risk exceeds $Qc$, then the continuation risk is smaller than the stopping risk, and hence, it is approximately optimal to stop at the first time such that $R_{n}^{\rm st}$ falls below $Qc$. This result, along with the approximation (55), yields $T^{0}\approx\min(\tau_{0},\tau_{1})$, where
@@
The test based of these stopping times is already optimal to second order. However, to make it third-order optimal one must choose the constant $Q$ carefully to account for the overshoots. Specifically, using this result, Lorden proves that the risks of an optimal rule and of the GSLRT are both connected to the risks of the family of one-sided tests $\tau_{a}(\theta)=\inf\left\{n:\lambda_{n}(\theta,\theta_{i})\geqslant a\right\}$, which are strictly optimal in the problem
$\rho(\theta,v)=\inf_{T}\left\{{\mathsf{E}}_{\theta}T+v{\mathsf{P}}_{\theta_{i}}(T<\infty)\right\}=\inf_{T}{\mathsf{E}}_{\theta}\left\{T+v\prod_{n=1}^{T}\frac{p_{\theta_{i}}(X_{n})}{p_{\theta}(X_{n})}\right\}$
if we set $a=\log[v{\mathcal{L}}(\theta,\theta_{i})]$. Therefore, if we now take $Q=1/{\mathcal{L}}(\theta,\theta_{i})$, then the resulting test will be nearly optimal to within $o(c)$. Since $\theta$ is unknown, we need to replace it with the estimate $\hat{\theta}_{n}$ to obtain
@@
In addition, it is interesting to compare Lorden’s approach with the Kiefer–Sacks test that stops the first time $R_{n}^{\rm st}$ becomes smaller than $c$. Lorden’s approach allows us to show that the test with the stopping time
@@
where ${\mathcal{L}}(\hat{\theta}_{n})={\mathcal{L}}(\hat{\theta}_{n},\theta_{1})$ if $\hat{\theta}_{n}<\theta^{*}$ and ${\mathcal{L}}(\hat{\theta}_{n})={\mathcal{L}}(\hat{\theta}_{n},\theta_{0})$ otherwise, is nearly optimal to within $o(c)$. Recall that the factor $\zeta(\theta,\theta_{i})=I(\theta,\theta_{i})^{-1}{\mathcal{L}}(\theta,\theta_{i})$ provides a necessary correction for the excess over the thresholds at stopping; see (51). This gives a considerable improvement over the Kiefer–Sacks test that ignores the overshoots, and not necessarily in the case of testing close hypotheses when ${\mathcal{L}}(\theta,\theta_{i})\ll 1$, but it may be important even if the parameter values are well separated. For example, in the binomial case with the success probabilities $\theta_{1}=0.6$ and $\theta_{0}=0.4$, ${\mathcal{L}}(\theta_{1},\theta_{0})\approx 1/15$, so Lorden’s test will stop much earlier.
Observe, however, that the implementation of Lorden’s fully optimized GSLRT may be problematic since usually computing the numbers $\zeta(\theta,\theta_{i})$ is not possible analytically except for some particular models such as the exponential. For example, when testing the mean in the Gaussian case, these numbers can be computed only numerically. Siegmund’s (1985) corrected Brownian motion approximations can be used but these are of sufficient accuracy only when the difference between $\theta$ and $\theta_{i}$ is relatively small. Therefore, for practical purposes, only partially optimized solutions, which provide $O(c)$-optimality, are always feasible. A work around is a discretization of the parameter space.

### v2

@@ 2.3 Lorden’s 2-SPRT and the Kiefer–Weiss Minimax Optimality
with $b(\theta)$ being a convex and infinitely differentiable function on $\widetilde{\Theta}\subset\Theta$, it is feasible to identify the nearly least favorable point $\theta^{*}(\alpha_{0},\alpha_{1},\theta_{0},\theta_{1})$ such that the $2$-SPRT with thresholds $a_{i}=\log(1/\alpha_{i})$ achieves second-order asymptotic minimaxity, meaning the residual term in the discrepancy between the expectation of the sample size of the optimal test and the 2-SPRT is of order $O(1)$ for small $\alpha_{i}$. Initially addressed by Huffman (1983), who proposed $\theta^{*}(\alpha_{0},\alpha_{1},\theta_{0},\theta_{1})$ leading to a residual term of order $o(|\log\alpha|^{1/2})$, this problem was further advanced by Dragalin and Novikov (1987), who demonstrated the second-order optimality of Huffman’s version of the 2-SPRT.
@@
As previously noted, the formulas $a_{i}=|\log\alpha_{i}|$, ensuring the inequalities $\alpha_{i}(\delta^{\star}(\theta))\leqslant\alpha_{i}$, tend to be overly conservative. A refinement can be achieved by observing that
@@
where, asymptotically as $a_{i}\to\infty$,
@@
For the non-arithmetic case, $\zeta_{i}^{\theta}$ can be computed using the renewal-theoretic argument similar to (11) and (13):
$\zeta_{i}^{\theta}=\frac{1}{I(\theta,\theta_{i})}\exp\left\{-\sum_{n=1}^{\infty}\frac{1}{n}\left[{\mathsf{P}}_{\theta}(\lambda_{i}^{\theta}(n)>0)+{\mathsf{P}}_{i}(\lambda_{i}^{\theta}(n)\leqslant 0)\right]\right\},$
where $I(\theta,\theta_{i})=(\theta-\theta_{i})\overset{\bm{.}}{b}(\theta)-(b(\theta)-b(\theta_{i}))$ are the Kullback–Leibler numbers.
The efficiency of Lorden’s 2-SPRT in stopping to both reject and accept the null was appealing in clinical trial designs, where these actions are known as efficacy and futility stopping. Lai and Shih (2004) extended the 2-SPRT from the fully sequential setting to the group sequential setting while maintaining its efficiency, and balancing the tradeoff between efficacy and futility stopping.
@@ 2.4 Near Uniform Optimality of the GLR SPRT for Composite Hypotheses
For practical purposes, it is considerably more significant to devise tests that minimize the expected sample size ${\mathsf{E}}_{\theta}[T]$ for all possible parameter values (i.e., uniformly optimal) rather than to address the minimax Kiefer–Weiss problem of minimizing ${\mathsf{E}}_{\theta}[T]$ at a least favorable point. In this section, our primary objective is to explore the design of sequential tests that are at least approximately uniformly optimal for small error probabilities or asymptotically Bayesian for a small cost of observations for testing composite hypotheses.
Consider a sequence of i.i.d. observations $X_{1},X_{2},\dots$ originating from a common distribution ${\mathsf{P}}_{\theta}$ with density $f_{\theta}(x)$ with respect to some non-degenerate sigma-finite measure, where the $\ell$-dimensional parameter $\theta=(\theta_{1},\dots,\theta_{\ell})$ belongs to a subset $\Theta$ of the Euclidean space $\mathbb{R}^{\ell}$. The parameter space $\Theta$ is partitioned into $3$ disjoint sets $\Theta_{0},\Theta_{1}$ and ${\mathsf{I}_{\mathrm{in}}}$, i.e., $\Theta=\Theta_{0}\,{\textstyle\bigcup}\,\Theta_{2}\,{\textstyle\bigcup}\,{\mathsf{I}_{\mathrm{in}}}$. The objective is to test the two composite hypotheses $\operatorname{\mathcal{H}}_{0}:\theta\in\Theta_{0}$ against $\operatorname{\mathcal{H}}_{1}:\theta\in\Theta_{1}$. The subset ${\mathsf{I}_{\mathrm{in}}}$ of $\Theta$ denotes an indifference zone where the loss $L(\theta,d)$ associated with correct or incorrect decisions $d$ is zero, i.e., no constraints on the probabilities ${\mathsf{P}}_{\theta}(d=i)$ are imposed if $\theta\in{\mathsf{I}_{\mathrm{in}}}$. The introduction of an indifference zone is typically motivated by the recognition that in many applications, the correct action is not crucial and often not even feasible when the hypotheses are very close. However, in principle ${\mathsf{I}_{\mathrm{in}}}$ may be an empty set.
We aim to find a sequential test $\delta=(T,d)$ that minimizes the expected sample size ${\mathsf{E}}_{\theta}[T]$ uniformly for all $\theta\in\Theta$ in the class of tests ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ in which the maximal error probabilities $\sup_{\theta\in\Theta_{i}}{\mathsf{P}}_{\theta}(d\neq i)$ are upper-bounded by the given values:
@@
Thus, we are interested in the frequentist problem of finding a test $\delta_{\rm opt}$ such that
@@
Kiefer and Sacks (1963) showed that the procedure $\widetilde{\delta}(c)=(\widetilde{T}(c),\widetilde{d}(c))$ with the stopping time $\widetilde{T}(c)=\inf\left\{n\geqslant 1:R_{n}^{\rm st}(S_{n})\leqslant c\right\},$ proposed by Schwarz (1962), is also first-order asymptotically Bayes. In other words, for any prior distribution $W$, $\rho_{c}^{W}(\widetilde{\delta}(c))$ behaves asymptotically like $\inf_{\delta}\rho_{c}^{W}(\delta)$ as $c\to 0$. Lorden (1967) refined this result by introducing the stopping region as the first $n$ such that $R_{n}^{\rm st}(S_{n})\leqslant Qc$, where $Q$ is a positive constant, and demonstrated that it can be made second-order asymptotically optimal, i.e., $\inf_{\delta}\rho_{c}^{W}(\delta)=\rho_{c}^{W}(\widetilde{\delta}(Qc))+O(c)$ as $c\to 0$, while $\inf_{\delta}\rho_{c}^{W}(\delta)=O(c|\log c|)$. It’s noteworthy that the problem addressed by Lorden (1967) is more general than what we are discussing here since it encompasses general i.i.d. models, not limited to exponential families, and multiple-decision cases. Additionally, see Lorden (1972) for multiple hypotheses in one-parameter exponential families.
A significant advancement in Bayesian theory for testing separated hypotheses about the parameter of the one-parameter exponential family (30) was made by Lorden (1977b) (an unpublished manuscript). In this work, Lorden demonstrated that the family of GSLRTs can be devised to ensure third-order asymptotic optimality. This implies that they achieve the Bayes risk to within $o(c)$ as $c\to 0$.
Lorden provided sufficient conditions for families of tests to be third-order asymptotically Bayes and presented examples of such procedures based not only on the Generalized Likelihood Ratio (GLR) approach but also on mixtures of likelihood ratios. Furthermore, the error probabilities of the GSLRTs were evaluated asymptotically as a consequence of a general theorem on boundary-crossing probabilities.
Due to the significance of this work, let’s delve into a more detailed overview of Lorden’s theory. It’s worth noting that the paper by Lorden (1977b) extends the results obtained by Lorden (1977a) for multiple discrete cases, which we discussed in Subsection 2.1.2, to the continuous parameter case.
@@
the LLR between points $\theta$ and $\theta_{i}$.
@@
Note the crucial difference between Schwarz’s GSLRT (35) and Lorden’s GSLRT (38). In the Schwarz test, $h_{i}\equiv 1$ and the threshold is set as $a=|\log c|$. However, in the Lorden test, two innovations emerge. Firstly, the threshold is reduced by $\tfrac{1}{2}\log|\log c|$, and secondly, adaptive weights $h_{i}(\hat{\theta}_{n})$ are incorporated into the GLR statistic. Since the stopping times $\widehat{T}_{i}$ can be written as
@@
Lorden’s GSLRT can alternatively be perceived as the GSLRT with curved adaptive boundaries
@@
which depend on the behavior of the MLE $\hat{\theta}_{n}$. These two innovations render this modification of the GLR test nearly optimal.
Given the complexity of Lorden’s formal mathematical proof, we offer a heuristic sketch that captures the main ideas of the approach. The Bayesian perspective naturally guides us toward the mixture LR statistics
@@
where $i=0$ if $\hat{\theta}_{n}\leqslant\theta^{*}$ and $i=1$ otherwise. These approximations stem from Laplace’s method for asymptotic integral expansions, and its variations.
Subsequently, Lorden demonstrated the existence of $Q>1$ such that if the stopping risk exceeds $Qc$, then the continuation risk becomes smaller than the stopping risk. Therefore, it is approximately optimal to stop at the first instance such that $R_{n}^{\rm st}$ falls below $Qc$. This finding, coupled with the approximation (43), results in $T_{\rm opt}\approx\min(\tau_{0},\tau_{1})$, where
@@
The test based on these stopping times is already optimal to the second order. However, to achieve third-order optimality, one must carefully choose the constant $Q$ to address the overshoots. Specifically, leveraging this result, Lorden demonstrates that the risks of an optimal rule and of the GSLRT are both linked to the risks of the family of one-sided tests $\tau_{a}(\theta)=\inf\{n:\lambda_{n}(\theta,\theta_{i})\geqslant a\}$, which are strictly optimal in the problem:
$\rho(\theta,v)=\inf_{T}\left\{{\mathsf{E}}_{\theta}T+v{\mathsf{P}}_{\theta_{i}}(T<\infty)\right\}=\inf_{T}{\mathsf{E}}_{\theta}\left\{T+v\prod_{n=1}^{T}\frac{p_{\theta_{i}}(X_{n})}{p_{\theta}(X_{n})}\right\}.$
If we set $a=\log[v{\mathcal{L}}(\theta,\theta_{i})]$, then by taking $Q=1/{\mathcal{L}}(\theta,\theta_{i})$ the resulting test will be nearly optimal to within $o(c)$. Since $\theta$ is unknown, we need to replace it with the estimate $\hat{\theta}_{n}$ to obtain
@@
It’s intriguing to compare Lorden’s approach with the Kiefer–Sacks test that stops the first time $R_{n}^{\rm st}$ becomes smaller than $c$. Lorden’s approach allows us to show that the test with the stopping time
@@
where ${\mathcal{L}}(\hat{\theta}_{n})={\mathcal{L}}(\hat{\theta}_{n},\theta_{1})$ if $\hat{\theta}_{n}<\theta^{*}$ and ${\mathcal{L}}(\hat{\theta}_{n})={\mathcal{L}}(\hat{\theta}_{n},\theta_{0})$ otherwise, is nearly optimal to within $o(c)$. It’s worth recalling that the factor $\zeta(\theta,\theta_{i})=I(\theta,\theta_{i})^{-1}{\mathcal{L}}(\theta,\theta_{i})$ provides a necessary correction for the excess over the thresholds at stopping; see (39). This offers a significant enhancement over the Kiefer–Sacks test, which disregards the overshoots. Notably, this improvement is not limited to testing close hypotheses when ${\mathcal{L}}(\theta,\theta_{i})\ll 1$. Even in cases where the parameter values are well-separated, this correction could be crucial. For instance, in the binomial case with the success probabilities $\theta_{1}=0.6$ and $\theta_{0}=0.4$, we have ${\mathcal{L}}(\theta_{1},\theta_{0})\approx 1/15$, so Lorden’s test will terminate much earlier.
Certainly, it’s important to note that implementing Lorden’s fully optimized GSLRT may encounter difficulties. This is primarily because computing the numbers $\zeta(\theta,\theta_{i})$ analytically is often not feasible, except for specific models such as the exponential. For instance, when testing the mean in the Gaussian case, these numbers can only be computed numerically. While Siegmund’s (1985) corrected Brownian motion approximations can be utilized, they are sufficiently accurate only when the difference between $\theta$ and $\theta_{i}$ is relatively small. Hence, for practical purposes, only partially optimized solutions, which provide $O(c)$-optimality, are typically feasible. A workaround involves discretizing the parameter space.